4. Halifax Manufacturing allows its customers to return merchandise for any reason up to 90 days...

4.

Halifax Manufacturing allows its customers to return merchandise for any reason up to 90 days after delivery and receive a credit to their accounts. All of Halifax's sales are for credit (no cash is collected at the time of sale). The company began 2021 with a refund liability of $340,000. During 2021, Halifax sold merchandise on account for $11,600,000. Halifax's merchandise costs it 75% of merchandise selling price. Also during the year, customers returned $575,000 in sales for credit, with $319,000 of those being returns of merchandise sold prior to 2021, and the rest being merchandise sold during 2021. Sales returns, estimated to be 5% of sales, are recorded as an adjusting entry at the end of the year.

Required:

1. Prepare entries to (a) record actual returns in 2021 of merchandise that was sold prior to 2021; (b) record actual returns in 2021 of merchandise that was sold during 2021; and (c) adjust the refund liability to its appropriate balance at year end.
2. What is the amount of the year-end refund liability after the adjusting entry is recorded?
Answer #1

ANSWER

Halifax Manufacturing
Opening refund liability         340,000 A
Less: Returns prior to 2021         319,000 B
Balance before adjustment           21,000 C=A-B
Cost of goods sold returned prior to 2021         239,250 D= B*75%
For 2021
Total Sales    11,600,000 E
Estimated Returns @ 5%         580,000 F=E*5%
Less: Already returned 256,000 G
(575,000 - 319,000 )
Refund liability to be created for 2021 324,000 H=F-G
Less: Balance already in the account            21,000 See C
Amount to be adjusted 303,000 I=H-C
Cost of goods sold returned in 2021
Sales returns in 2021         256,000 See G
Merchandise costs at 75%         192,000 J=G*75%
Inventory- estimated returns for 2021
Amount to be adjusted in refund liability         303,000 See I
Merchandise costs at 75%         227,250 K=I*75%
Answer 1
Journal Entries
Account Debit $ Credit $
Refund Liability         319,000 See Note B
Accounts Receivable 319,000
Inventory         239,250 See Note D
Inventory- estimated returns 239,250
Sales returns         256,000 See Note G
Accounts Receivable 256,000
Inventory         192,000 See Note J
Cost of goods sold 192,000
Sales returns         303,000 See Note I
Refund liability 303,000
Inventory- estimated returns         227,250 See Note K
Cost of goods sold 227,250
Answer 2- See Note H
The amount of the year end refund liability after the adjustment entry is recorded is- $ 324,000.

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************